ID |
Date |
Author |
Type |
Category |
Subject |
15806
|
Fri Feb 12 15:03:48 2021 |
Jordan | Update | SUS | CoM on 3"->2" Adapter Ring for SOS |
As it currently stands the Center of Mass of the Adapter Ring/Optic assembly is 0.0175" out of the plane formed by the suspension wire. See Attachments. The side plate, along with the EQ stops are hidden to show the CoM and the plane.
Note: The changes discussed in the meeting with Calum have not been added and are a work in progress. These changes include:
- Adding a 45 deg chamfer to the both parallel faces of the adapter ring. This along with a modified bracket for the EQ stops will allow for easier adjustment of the screws.
- Potentially changing material of adapter ring to stainless stell to more accurately emulate the mass of a 3" optic.
- Different adjustment mechanism of the "dumbell" at bottom of adapter ring to something similar to the VOPO suspension (will need to consult Calum further) |
Attachment 1: Screenshot_(1).png
|
|
Attachment 2: Screenshot_(3).png
|
|
Attachment 3: CoM.PNG
|
|
15805
|
Thu Feb 11 18:21:39 2021 |
gautam | Update | SUS | MC suspension glitches |
MC1 suspension is glitching again, so this is a good chance to install the new sat box and test it in the field. |
15804
|
Thu Feb 11 16:58:52 2021 |
rana | Summary | BHD | Satellite Amplifier Very Low frequency noise After modifications |
I expect that a single OSEM channel can't be better than 1e-10 m/rHz above 5 Hz, so probably something wrong in the calibration. 1.6 V/mm seems right to me, so could be some place else. |
15803
|
Thu Feb 11 11:10:05 2021 |
Anchal | Summary | BHD | Satellite Amplifier Very Low frequency noise After modifications |
Here is a proper measurement for PD transimpedance amplifier circuit in the Satellite amplifier box D1002818 S2100029. The input from rear DB25 connector was left open and measurement was taken with AC coupling with correction by the AC coupling transfer function (Zero at 0, pole at 160 mHz). I have calculated the input referred displacement noise by calculating the conversion factor of OSEM in A/m. From 40m/12470, old conversion factor of OSEM to output of sat amplifier was 1.6 V/mm. then, the transimpedance was 39.2 kOhm, so that must mean a conversion factor of 1.6e3/39.2 A/m. This I scaled with increased drive current by factor of 35/25 as mentioned in this document. The final conversion factor turned out to be around 57 mA / m. If someone finds error in this, please let me know.
There is excess noise in the low-frequency region below 5-6 Hz. If people think I should make a measurement of amplified noise to go further away from the instrument noise floor, let me know. |
Attachment 1: AfterChangesSpectrum_AC.zip
|
Attachment 2: D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf
|
|
Attachment 3: D1002818_S2100029_InputRefCurrentNoiseSpecAfterChanges.pdf
|
|
Attachment 4: D1002818_S2100029_InputRefDispNoiseSpecAfterChanges.pdf
|
|
15802
|
Wed Feb 10 21:14:03 2021 |
gautam | Update | Electronics | Production version of the HV coil driver tested |
Summary:
I did what I consider to be a comprehensive set of tests on the production version of the high voltage coil driver circuit. I think the performance is now satisfactory and the circuit is ready for the production build. Barring objections from anyone, I will ask Chub to place an order for components to stuff the 4 necessary units + 1 spare on Friday, 12 Feb (so that people have a full day to comment). A big thanks to Chub and the folks at JLCPCB for dealing with my incessant order requests and patiently supporting this build and letting me turn this around in 10 days - hopefully this is the end of this particular saga.
Schematic is here. All references to component designations are for v4 of the schematic.
Important design changes:
- All I/O to this board will be via D9 connectors. This will allow bypassing the high voltage stage in future suspensions while retaining the same cable config in the suspension drive, if that is desired. Some re-arrangement of the grouping of coils was also done for consistency with the planned upgrade.
- Differential receiving for the input signal from the Acromag. The nominal quad opamp is LT1125 but if we find noise issues (which I didn't), the OP497 has compatible PCB footprint.
- Added input protection dual diode D6 to protect the PA95 as recommended in the datasheet. This should protect the IC if (for example) the HV line isn't plugged in but the Acromag input is non-zero.
- Increased the feedback resistance from 30kohms to 12kohms. R16 through R21 are now 20k, while the old design had 5k. The purpose is to reduce the current demand in the feedback path, hopefully this opens up the number of DCPS we can use. To keep the overall gain of 31, the resistor R15 was upped from 1kohms to 4kohms.
- Feedback capacitance reduced from 15 uF to 3 uF. This was largely for space saving / ease of layout on the PCB. The resulting corner frequency is increased slightly from 0.35 Hz to 0.45 Hz but this doesn't have any imapct on the performance of the circuit at frequencies of interest (1/2/pi/R/C had R=30k, C=15uF, now R=120k, C=3uF).
- Added an R-C-R network at the output of the PA95, before the fast actuation signal is summed and sent to the OSEM coil.
- This is probably the most important change, noise-performance wise.
- The purpose of the network is to passively filter out the excess noise we saw at ~100 Hz (the corner from the 4kohm resistor + 3uF cap is at ~13 Hz, so factor of 10 filtering at 100 Hz, which was deemed sufficient, see earlier elogs in the thread).
- The Johnson noise contribution of the 20 kohm resistor is slightly higher than the original design which had a 25 kohm series resistor (but no R-C-R passive filter at the output of the PA95). But once again, this was deemed to have negligible effect on the performance at frequencies of interest to us.
- The total current driving capability of the circuit is almost unchanged since the 20kohm + 4kohm nearly equals the old 25kohm resistance.
- Made the Vmon paths for monitoring the HV output of the PA95 differential sending, seems like a good practise to follow for all new designs.
- Added on-board bypass capacitors (2 x 10uF WIMA film caps) for cleaning up the HV supply noise.
Tests:
A series of tests were done. Note that only 1 channel was stuffed (I am out of PA95s), and the HP power supplies borrowed from Rich were used for the HV rails. For the +/-18V, a regular bench-top unit was used.
- Low voltage stage tests
- TF of the differential receiving stage was measured and the overall unity gain and corner at 24kHz were verified, see Attachment #1.
- With the input of the circuit grounded, I measured the noise of the circuit at various points (see legends on Attachment #2). I didn't bother to do a detailed verification against a SPICE model as the levels seemed roughly what is expected.
- Overall noise performance with HV stage enabled
- For a range of drive currents, generated by applying the appropriate voltage using an Acromag XT1541 DAC module to the J1 connector, I measured the voltage on the circuit side of the 20 kohm resistor (by clipping onto the resistor leg. Note that the path to ground for the current was provided by connecting a 20 ohm resistor between pins 1 and 6 on J3a, and then grounding pin 6.
- Results are shown in Attachment #3.
- For the drive currents at the edge of the range of operation, there is a small excess relative to lower drive currents. My best hypothesis for why this is happening is that the HV rail is too close to the requested output voltage (the HP units are rated for 320V, which is borderline if we want 300V at the output of the PA95). In any case, the R-C-R passive filter means that above ~60 Hz, there is excellent agreement between model and measurement.
- Time domain tests
- Used a function generator. to drive a 50 mHz, 3Vpp sine wave to the "Bias Input" (=J1), and monitored (i) pickoff of drive signal, (ii) High voltage output at the circuit side of the 20kohm resistor, and (iii) the Vmon output (=pins 1/6 on J4), all on a 100 MHz Tektronix scope.
- Results shown in Attachment #4. Once again, I see no red flags.
- While I had the unit hooked up to the scope, I also checked the time domain signal with the scope set to 100 ns/div time base. I saw no evidence of any oscillatory features, either when no input signal was applied, or when a DC signal was provided (in which case the scope was set to AC coupling). 👍
- Checked that the protection diodes at various locations in the circuit work.
- Checked the pin-mapping on all 6 D9 connectors is consistent with the top level diagram in the schematic.
PCB remarks:
As I was stuffing the board, I noticed a few improvements that can be made. Just noting these here for documentation purposes - these changes are mostly aesthetic and I personally see no need to order another set of PCBs.
- In some places, the silkscreen designators don't have the correct "orientation" relative to the component they are designating. I didn't find any serious ambiguity in terms of being misled to stuff the wrong component onto the wrong pads, but in the spirit of doing a professional job...
- The current limiting resistors on the +/-430V LEDs (R37/R38) have footprints for leaded components rather than SMT (which is what the +/-15V LEDs have).
- R45 and R46, the current limiting resistors for the rear panel power indicator LEDs, have 0805 footprint rather than 1206.
- When I drew up the PCB, R23, the 4kohm resistor in the R-C-R network, was set up as a 1W resistor. Let's say there can be 15 mA flowing through this resistor - the power dissipated is 15e-3 ^2 * 4e3 is 0.9W, which is uncomfortably close to the limit. For all the tests above, I used a 3W resistor, and didn't find any serious noise issues. The drilled holes are a little tight for this higher power rated resistor, but I don't think this is a showstopper.
Communications with Apex:
I've been talking to support at Apex, and pointed out that I couldn't match the SPICE model performance even for a simple non-inverting amplifier with the PA95. The feedback I got from them was that
- They don't optimize the SPICE models for input noise and so it was a nice coincidence that model and measurement are somewhat close (but not exactly).
- They recommend the PA194, which is actually advertised as "low-noise". The PA95 is apparently not a "low-noise" part, with its 2uVrms input noise.
Whiel the PA194 is compatible with our voltage and current requirements for this application, it is ~3x the cost, and seems like the R-C-R output filter allows us to realize the goal of 1pA/rtHz, so I'm inclined to stick with the PA95.
Production assembly:
I'd prefer to get as much of the board stuffed by Screaming Circuits as possible. It took me ~3 hours to stuff 1 channel + the power supply parts, standoffs etc. So I estimate it'll take me ~6 hours to stuff the entire board. So not the end of the world if we have to do it in-house. |
Attachment 1: inputDiffRecTF.pdf
|
|
Attachment 2: LVnoises.pdf
|
|
Attachment 3: totalNoise.pdf
|
|
Attachment 4: timeDomainTests.pdf
|
|
15801
|
Wed Feb 10 17:18:03 2021 |
Koji | Summary | BHD | Satellite Amplifier Output Offset measurements |
Testing the satellite amp i.e. PD driver
- To test the noise of the PD transimpedance amps: Leave the PD input open (do not short the terminal goes to the PD)
- To test the current noise of the LED drivers: Short the output with an appropriate Rs to have the nominal current.
- To test the overall noise level together with the LED/PD pair: Connect the dummy OSEM module.
Testing the coil drivers
- Short the output with an appropriate Rs. |
15800
|
Wed Feb 10 15:25:45 2021 |
gautam | Summary | BHD | Satellite Amplifier Output Offset measurements |
Why not just do this test with the dummy suspension box and CDS system? I think Rich's claim was that the intrinsic LED RIN was dominant over any drive current noise but we can at least measure the quadrature sum of the two (which is after all the relevant quantity in terms of what performance we can realize) and compare to a model. |
15799
|
Wed Feb 10 15:07:50 2021 |
Anchal | Summary | BHD | Satellite Amplifier Output Offset measurements |
I measured the output DC voltage of the satellite amplifier box at PDMon port when the PDA input was shorted and got following offsets:
CH |
Output Offset (mV) |
CH |
Output Offset (mV) |
1 |
6 |
5 |
750 |
2 |
140 |
6 |
120 |
3 |
350 |
7 |
537 |
4 |
40 |
8 |
670 |
However, I think I'm making a mistake while measuring this offset as well as all the noise measurements of this satellite amplifier box so far. Since it is a current input, transimpedance circuit, the noise of the circuit should be measured with open input, not closed. Infact, by shorting the PDA input, I'm giving DC path to input bias current of AD833 transimpedance amplifier to create this huge DC offset. This won't be the case when a photodiode is connected at the input which is a capacitor and hence no DC path is allowed. So my issue of offset was bogus and past two noise measurements in 40m/15797 and 40m/15793 are wrong. |
15798
|
Wed Feb 10 14:14:58 2021 |
gautam | Update | Electronics | Custom cables received |
We received the custom cables to test the new suspension electronics. They are under my desk. So we are ready.
This batch was a small one - the company says that they can make molded cables if we have a minimum order, something to consider I gues.s.
Update 1900 11 Feb: I verified that the pin outs of the cables are as we intended (for one set of each type of cable). Because this was a small order, the connectors have metal shells, and so for cable #2 (sat box to flange), the two shells are shorted to each other. I can't verify if the shield is isolated from the shell on J5 without cutting open the cable. One thing that occurred to me is that we should give pins 5,8,11 on J4 and 16,20,24 on J5 (respectively) unique identifiers. They should only be shorted to GND on the circuit board itself. To be fixed for the next iteration. I uploaded some photos here.
I was unable to measure the capacitance of the cable using the LCR meter, and didn't opt to try any other method. |
Attachment 1: satWiring.pdf
|
|
15797
|
Wed Feb 10 11:45:59 2021 |
Anchal | Summary | BHD | Satellite Amplifier Very Low frequency noise After modifications |
As suggested, I wrapped the satellite amplifier box D10028128 S2100029 in blanket and foam and took very low frequency spectrum starting from 32 mHz to 3 Hz. The results are attached along with stiched high frequency measurements from 40m/15793.
Very Low Frequency Spectrum Measurement
- D1002818 S2100029 box was powered and covered in a foam blanket.
- Additionally, it was covered from all sides with foam to reduce wind and temperature effects on it.
- The rear panel DB25 connector was connected to a breakout board where pins od PDA input and GND were shorted, shorting the transimpedance circuit input.
- The output was read from PDMon DB9 output at front panel which was converted to 4 BNC channels using breakout board.
- Two channel noise was measured at once using D1002818_SP.yml parameter file.
- Instrument noise at all the used input ranges were measured separately by shorting the input of the BNC cables.
Edit Wed Feb 10 15:14:13 2021 :
THIS MEASUREMENT WAS WRONG. SEE 40m/15799. |
Attachment 1: FrontsideLook.jpg
|
|
Attachment 2: BacksideLook.jpg
|
|
Attachment 3: InnerFoamBlanket.jpg
|
|
Attachment 4: D1002818_S2100029_OutputNoiseLFSpecAfterChanges.pdf
|
|
Attachment 5: D1002818_S2100029_InputRefCurrentNoiseLFSpecAfterChanges.pdf
|
|
Attachment 6: AfterChangesLFSpectrum.zip
|
15796
|
Thu Feb 4 15:14:55 2021 |
Yehonathan | Update | BHD | SOS assembly |
I gathered all the components I could find from the SOS towers and the cleanroom and put it all on the table next to the flow bench (See attachment).
I combed through the cleanroom cabinet for SOS parts but didn't find all the parts listed in the procurement spreadsheet. I did find some extra items that were not listed.
This table compares the quantities in the spreadsheet to the quantities collected on the table. Green rows are items I found more than in the procurement spreedsheet while red rows are items I found less.
ITEM |
DCC # |
Qty required |
Qty in procurement spreadsheet |
How much I found |
SENSOR/ACTUATOR PLATE |
D960002 |
14 |
21 |
21 |
SUSPENSION BLOCK |
D960003 |
7 |
9 |
9 |
TOWER BASE |
D960004 |
7 |
10 |
11 |
RIGHT SIDE PLATE |
D960005 |
7 |
12 |
13 |
LEFT SIDE PLATE |
D960006 |
7 |
12 |
13 |
UPPER MIRROR CLAMP |
D960007 |
7 |
8 |
7+1 teflon
|
LOWER CLAMP |
D960008-1 |
7 |
8 |
8 |
LOWER CLAMP, OPPOSITE |
D960008-2 |
7 |
8 |
8 |
WIRE CLAMP |
1205308-1 |
10 |
17 |
9 |
CLAMP, SUSPENSION BLOCK |
D960134 |
14 |
19 |
21 |
STIFFENER PLATE |
D960009 |
7 |
9 |
9 |
DUMBBELL STANDOFF |
D970075 |
50 |
10 |
7 |
SAFETY STOP, LONG |
D970313 |
14 |
2 |
10 |
OSEM assy |
D960011 |
35 |
2 |
13 wire wound osem housings (gold) |
WIRE STANDOFF |
D970187 |
20 |
7 |
0 |
GUIDE ROD |
D970188 |
10 |
9 |
0 |
MAGNET |
D960501 |
50 |
54 |
51 rusted + 37 slightly rusted. Didn't put on table |
SAFETY STOP, SMALL |
D970312 |
28 |
0 |
4 |
SAFETY STOP |
D970311 |
28 |
0 |
16+9 stained w/o spring
|
SS Spring Plunger |
8498A999 |
35 |
4 |
27 |
|
Attachment 1: 20210204_144007.jpg
|
|
15795
|
Wed Feb 3 21:28:02 2021 |
gautam | Update | CDS | CDS crash and CDS/IFO recovery |
I am just reporting my experience - this may be a new failure mode but I don't think so. In the new RTCDS, the ntp server for the FEs are the FB, to which they are synced by timesyncd. The FB machine itself has the ntpd service installed, and so is capable of synching to an NTP server over the internet, but also serving as an NTP server for the FEs. The timesyncd daemon may not have started correctly, or the NTP serving from FB got interrupted (for example), but that's all just speculation. |
15794
|
Wed Feb 3 18:53:31 2021 |
Koji | Update | CDS | CDS crash and CDS/IFO recovery |
Really!? I didn't reboot the machines between "sudo date" and "rtcds start c1x0*". I tried rtcds. If it didn't work, it used date. Then tried rtcds. (repeat) The time was not synched at all wrt the time zones and also the time. There were 1~3 sec offset besides the TZ problem.
|
15793
|
Wed Feb 3 16:27:19 2021 |
Anchal | Summary | BHD | Satellite Amplifier Transfer Functions and noise After modifications |
I have made modifications recommended in this doc. The changes made are:
- R24: 19.6k to 4.99k Ohms
- R20: 19.6k to 4.99k Ohms
- R23: 787 to 499 Ohms
- Removed C16.
I took transfer function measurements, fitted them with zeros and poles and plotted it against the zero model of the circuit. The zeros and poles we intended to shift are matching well with 3Hz zero and 30 Hz pole. The later pole at 1500 Hz is at a higher value from what is predicted by zero.
I also took noise measurements and they are in good agreement with the noise predicted by zero.
Edit Wed Feb 10 15:14:13 2021 :
THE NOISE MEASUREMENT WAS WRONG HERE. SEE 40m/15799. |
Attachment 1: D1002818_S2100029_TFAfterChanges.pdf
|
|
Attachment 2: D1002818_S2100029_OutputNoiseSpecAfterChanges.pdf
|
|
Attachment 3: D1002818_S2100029_InputRefferedNoiseSpecAfterChanges.pdf
|
|
Attachment 4: D1002812_S2100029_After_Modifications_Feb3.jpg
|
|
Attachment 5: AfterChanges.zip
|
15792
|
Wed Feb 3 15:24:52 2021 |
gautam | Update | CDS | CDS crash and CDS/IFO recovery |
Didn't get a chance to comment during the meeting - This was almost certainly a coincidence. I have never had to do this - I assert, based on the ~10 labwide reboots I have had to do in the last two years, that whether the timing errors persist on reboot or not is not deterministic. But this is beyond my level of CDS knowledge and so I'm happy for Rolf / Jamie to comment. I use the reboot script - if that doesn't work, I use it again until the systems come back without any errors.
Quote: |
This looked like the usual timing issue. It looked like "ntpdate" is not available in the new system. (When was it updated?)
The hardware clock (RTC) of these hosts are set to be PST while the functional end host showed UTC. So I copied the time of the UTC time from the end to the vertex machines.
For the time adjustment, the standard "date" command was used
> sudo date -s "2021-02-03 07:11:30"
This made the trick. Once IOP was restarted, the "DC" indicators returned to **Green**, restarting the other processes were straight forward and now the CDS indicators are all green.
|
I don't think this is a problem, the NTP synchronization is handled by timesyncd now.
Quote: |
NTP synchronization is not active. Is this OK?
|
I defer restoring the LSC settings etc since I guess there is not expected to be any interferometer activity for a while. |
15791
|
Tue Feb 2 23:29:35 2021 |
Koji | Update | CDS | CDS crash and CDS/IFO recovery |
I worked around the racks and the feedthru flanges this afternoon and evening. This inevitably crashed c1lsc real-time process.
Rebooting c1lsc caused multiple crashes (as usual) and I had to hard reboot c1lsc/c1sus/c1ioo
This made the "DC" indicator of the IOPs for these hosts **RED**.
This looked like the usual timing issue. It looked like "ntpdate" is not available in the new system. (When was it updated?)
The hardware clock (RTC) of these hosts are set to be PST while the functional end host showed UTC. So I copied the time of the UTC time from the end to the vertex machines.
For the time adjustment, the standard "date" command was used
> sudo date -s "2021-02-03 07:11:30"
This made the trick. Once IOP was restarted, the "DC" indicators returned to **Green**, restarting the other processes were straight forward and now the CDS indicators are all green.
controls@ c1iscex :~ 0$ timedatectl
Local time: Wed 2021-02-03 07:35:12 UTC
Universal time: Wed 2021-02-03 07:35:12 UTC
RTC time: Wed 2021-02-03 07:35:26
Time zone: Etc/UTC (UTC, +0000)
NTP enabled: yes
NTP synchronized: no
RTC in local TZ: no
DST active: n/a
NTP synchronization is not active. Is this OK?
With the recovered CDS, the IMC was immediately locked and the autolocker started to function after a few pokes (like manually running of the "mcup" script). However, I didn't see any light on the AS/REF cameras as well as the test mass faces. I'm sure the IMC alignment is OK. This means the TTs are not well aligned.
So, burtrestored c1assepics with 12:19 snapshot. This immediately brought the spots on the REFL/AS.
Then the arm were aligned, locked, and ASSed. I tried to lock the FP arms. The transmissions were at the level of 0.1~0.3. So some manual alignment of ITMY and BS were necessary. After having the TRs of ~0.8, I still could not lock the arms. The signs of the servo gains were flipped to -0.143 for X arm and -0.012 for Y arm, and the arms were locked. ASS worked well and the ASS offsets were offloaded to the SUSs.
|
15790
|
Tue Feb 2 18:24:54 2021 |
Koji | Update | BHD | SOS assembly |
You can remove the components of the optical table enclosure (black ones) and use the optical table as your working area too.
|
15788
|
Tue Feb 2 17:09:17 2021 |
yehonathan | Update | BHD | SOS assembly |
I set up a working area on the table next to the south flow bench (see attachment). I also brought in a rolling table for some extra space.
I covered all the working surfaces with a foil from the big roll between 1x3 and 1x4.
I took the SOSs, SOS parts and the OSEMS from the MC2 table to the working area.
I cleaned some LN Allen keys with isopropanol and put them on the working table, please don't take them. |
Attachment 1: 20210202_165501.jpg
|
|
Attachment 2: 20210202_162452.jpg
|
|
15787
|
Tue Feb 2 11:57:46 2021 |
Anchal | Summary | BHD | HAM-A Coil Driver measurements After modifications TF and Noise S2100028 |
I have made the modifications on the other board D1100687 S2100028 as well. The measurements were taken as mentioned in 40m/15784. All conclusions remain the same as 40m/15784. The attached zip file contains all measurement data, before and after the modifications.
Edit Wed Feb 3 16:44:51 2021 :
Added zero modeled noise in the noise spectrum curves. The acquisition mode curves are in agreement with the model. The noise in Run mode is weirdly lower than predicted by zero. |
Attachment 1: D1100687_S2100028_After_Modifications_Feb01_2021.jpg
|
|
Attachment 2: D1100117_S2100028_TF.pdf
|
|
Attachment 3: D1100117_S2100028_Voltage_Noise_Spectrum.pdf
|
|
Attachment 4: D1100117_S2100028_Current_Noise_Spectrum.pdf
|
|
Attachment 5: AfterChanges.zip
|
15786
|
Mon Feb 1 12:30:21 2021 |
gautam | Update | Electronics | More careful characterization |
Summary:
- Swapping out the KEPCO HV supplies (linear) I was using for a pair of HP6209s I borrowed from Rich has improved the noise performance somewhat.
- However, there is still an excess relative to the model. I confirmed that this excess originates from the PA95 part of the circuit (see details).
- The bypass capacitors don't seem to have any effect on the measured ripple from these HP6209s. Maybe they're internally fitted with some 10uF or similar bypass caps?
- The production version of this board, with several improvements (after discussions with Koji and Rich), are on the DCC. They're being fabbed right now and will arrive in ~1 week for more bench testing.
Power supply bypassing [updated 10pm]:
As mentioned earlier in this thread, I prepared a box with two 10uF, 1kV rated capacitors to bypass the high-voltage rails (see inset in the plot), to see if that improves the performance. However, in measuring the voltage ripple directly with the SR785 (no load connected), I don't see any significant difference whether the decoupling caps are connected or not, see Attachment #1. For this, and all other HV measurements made, I used this box to protect the SR785. One hypothesis is that this box itself is somehow introducting the excess noise, maybe because of leakage currents of the diode pair going into the 1Mohm SR785 input impedance, but I can't find any spec for this, and anyway, these diodes should be at ground potential once the transient has settled and the DC blocking capacitor has charged to its final value.
Note that the 10uF caps have an ESR of 7.2 mOhms. The HP6209 has a source impedance "<20mOhm" when operated as a CV source, per the datasheet. So perhaps this isn't so surprising? The same datasheet suggests the source impedance is 500 mOhms from 1kHz to 100 kHz, so we should see some improvement there, but I only measured out to 2 kHz, and I didn't take much effort to reduce these crazy peaks so maybe they are polluting the measurement out there. There must also be some continuous change of impedance, it cannot be <20 mOhm until 1 kHz and then suddenly increase to 500 mOhms. Anyways, for this particular circuit, the nosie DC-1kHz is what is important so I don't see a need to beat this horse more.
Simplified circuit testing:
I decided to see if I can recover the spec'd voltage noise curve from the PA95 datasheet. For this, I configured the PA95 as a simple G=31 non-inverting amplifier (by not stuffing the 15 uF capacitor in the feedback path). Then, with the input grounded, I measured the output voltage noise on the circuit side of the 25kohm resistor (see inset in Attachment #2). To be consistent, I used the DC blocking box for this measurement as well, even though the output of the PA95 under these test conditions is 0V. Once again, there is considerable excess around ~100 Hz relative to a SPICE model. On the basis of this test, I think it is fair to say that the problem is with the PA95 itself. As far as I can tell, I am doing everything by the book, in terms of having gain > 10, using a sufficiently large compensaiton cap, HV rail decoupling etc etc. Note that the PA95 is a FET input opamp, so the effects of input current noise should be negligible. The datasheet doesn't provide the frequency dependence, but if this is just shot noise of the 1200 pA input bias current (for 300 V rails, per the spec), this is totally negligible, as confirmed by LTspice.
In the spirit of going step-by-step, I then added the feedback capacitor, and still, measured noise in excess of what I would expect from my model + SR785 measurement noise.
Integrated circuit testing:
After the above simplified test, I stuffed a full channel as designed, and tested the noise for various drive currents. To best simulate the operating conditions, an Acromag XT1541 was used to set the DC voltage that determines the drive current through the 25 kohm resistor. The measurements were made on the circuit side of this resistor (I connected a 20ohm resistor to ground to simulate the OSEM). As shown in Attachment #3, the noise with these HP6209 supplies is significantly better than what I saw with the KEPCO supplies, lending further credence to the hypothesis that insufficient PSRR is the root of the problem here. I've added subplots in a few different units - to be honest, I think that reaching this level of measured displacement noise at the 40m at 100 Hz would already be pretty impressive.
So what's next?
The main design change is that a passive R-C-R (4k-3uF-20k) replaces the single 25kohm resistor at the output of the PA95.
- This allows similar current drive range.
- But adds an LPF to filter out the observerd excess noise at 100 Hz.
Let's see if this fixes the issue. Not that I've also added a pair of input protection diodes to the input of the PA95 in the new design. The idea is that this would protect the (expensive) PA95 IC from, for example, the unit being powered with the +/- 18V rail but not the +/- 300 V rail. As I type this, however, I wonder if the leakage current noise of these diodes would be a problem. Once again, the datasheet doesn't provide any frequency dependence, but if it's just the shot noise of the 1nA expected when the diodes are not reverse biased (which is the case when the PA95 is operating normally since both inputs are at nearly the same potential), the level is ~20 fA/rtHz, comparable to the input current noise of the PA95, so not expected to be an issue. In the worst case, the PCB layout allows for this component to just be omitted. |
Attachment 1: HVPS.pdf
|
|
Attachment 2: HV_testckt.pdf
|
|
Attachment 3: totalNoise.pdf
|
|
15785
|
Fri Jan 29 17:57:17 2021 |
Anchal | HowTo | CDS | Acromag wiring investigation |
I found a white paper from Acromag which discusses how to read differential signal using Acromag units. The document categorically says that differential signals are always supposed to be transmitted in three wires. I provides the two options of either using the RTN to connect to the signal ground (as done in Attachment 3) or locally place 10k-100k resistors between return and IN+ and IN- both (Attachment 2).
I have provided possible scenarios for these.
Using two wires to carry differential signal (Attachment 1):
- I assume this is our preferential way to connect.
- We can also assume all single-ended inputs as differential and do a signal condition agnostic wiring.
- Attachment 3 show what were the results for different values of resistors when a 2Hz 0.5V amplitude signal from SR785 which as converted to differential signal using D1900068 was measured by acromag.
- The connection to RTN is symmetrical for both inputs.
Using three wires to carry differential signal (Attachment 2):
- This is recommended method by the document in which it asks to carry the GND from signal source and connect it to RTN.
- If we use this, we'll have to be very cautious on what GND has been shorted through the acromag RTN terminals.
- This would probably create a lot of opportunities for ground loops to form.
Using an acromag card without making any connection with RTN is basically not allowed as per this document. |
Attachment 1: GeneralLabWiringDiffWith2Wires.pdf
|
|
Attachment 2: GeneralLabWiringDiffWith3Wires.pdf
|
|
Attachment 3: DiffReadResistorbtwnINandRTN.pdf
|
|
15784
|
Fri Jan 29 15:39:30 2021 |
Anchal | Summary | BHD | HAM-A Coil Driver measurements After modifications TF and Noise S2100027 |
I fitted zeros and poles in the measured transfer function of D1100687 S2100027 and got zeros at 130 Hz and 234 Hz and poles at 10Hz and 2845 Hz. These values are different from the aimed values in this doc, particularly the 234Hz zero which was aimed at 530 Hz in the doc.
I also took the noise measurement using the same method as described in 40m/15780. The noise in Acquisition mode seems to have gone up in 10 Hz - 500 Hz region compared to the measurement in 40m/15780 before the modifications.
All channels are consistent with each other.
Edit Mon Feb 1 12:24:14 2021:
Added zero model prediction after the changes. The measurements match with the predictions.
Edit Wed Feb 3 16:46:59 2021:
Added zero modeled noise in the noise spectrum curves. The acquisition mode curves are in agreement with the model. The noise in Run mode is weirdly lower than predicted by zero. |
Attachment 1: D1100687_S2100027_After_Modifications_Jan28.jpg
|
|
Attachment 2: D1100117_S2100027_TF.pdf
|
|
Attachment 3: D1100117_S2100027_Voltage_Noise_Spectrum.pdf
|
|
Attachment 4: D1100117_S2100027_Current_Noise_Spectrum.pdf
|
|
Attachment 5: AfterChanges.zip
|
15783
|
Thu Jan 28 22:34:21 2021 |
gautam | Update | SUS | De-whitening |
Summary:
- We will need de-whitening filters for the BHD relay optics in order to meet the displacement noise requirements set out in the DRD. I think these need not be remotely switchable (depends on specifics of LO phase control scheme). SR2, PR2 and PR3 can also have the same config, and probably MC1, MC3 as well.
- We will need de-whitening filters for the non test mass core IFO optics (PRM, SRM, BS, and probably MC2).
- I am pretty sure we will not be able to have sufficient DAC range for the latter class of optics if we have to:
- Supply the DC bias.
- Do the LSC and ASC actuation in the presence of reasonable sensing noise levels.
- Engage de-whitening to low-pass-filter the DAC noise at ~200 Hz.
Details:
Attachment #1 shows the DAC noise models for the General Standards 16-bit and 18-bit DACs we are expecting to have.
- The 16-bit model has been validated by me at the 40m a few years ago.
- We have never used the 18-bit flavor at the 40m, and there are all manner of quirks apparently related to zero crossings and such. So the noise may be up to x2 higher (we won't have as much freedom necessarily as the sites to bias the DAC on one side of the zero crossing if we also need to use the same DAC channel to supply the DC bias current for alignment.
Attachment #2 shows the expected actuation range for DC optic alignment, assuming we use the entire DAC range for this purpose.
- Clearly, we need to do other things with the same DAC channels as well, so this is very much an upper bound of what will be possible.
- Let's assume we will not go lower than 100ohms.
- For all new optics we are suspending, we should aim to get the pitch balancing to within 500urad. With a 2x2m=4m optical lever arm, this corresponds to a 2mm spot shift. Should be doable.
- This could turn out to be a serious problem for PRM, BS and SRM if we hope to measure squeezing - the <AUX DOF>-->DARM coupling could be at the level of -40dB, and at 200 Hz, the DAC noise would result in PRCL/MICH/SRCL noise at the level of ~10^-15m/rtHz, which would be 10^-17m/rtHz in DARM. I don't think we can get 20dB of feedforward cancellation at these frequencies. For demonstrating locking using a BHD error signal, maybe this is not a big deal.
Attachment #3 shows the current and proposed (by me, just a rough first pass, not optimized in any way yet) de-whitening filter shapes. These shapes can be tweaked for sure.
- The existing de-whitening filter is way too aggressive. FWIW, the DRD "models" a "4th order Chebyshev low pass filter" which doesn't exist anywhere as far as I know.
- Since the DAC noise is below 1 uV/rtHz at all frequencies of interest, we never need to have >60dB de-whitening anywhere as the input referred noise of any circuit we build will exceed 1 nV/rtHz.
- I propose 3 poles, 3 zeros. In the plot, these poles are located at 30Hz, 50Hz, 2kHz, and the zeros are at 300 Hz, 300 Hz, 800 Hz.
- The de-whitening is less agressive below 100 Hz, where we still need significant LSC actuation ability. Considering the sensing noise levels at the 40m, I don't know if we can have reasonable LSC and ASC loop shapes and still have the de-whitening.
- Once again, PRM, SRM and BS will be the most challenging.
- For the BHD relay optics, once we have the de-whitening, we won't have the option of turning on a high-frequency (~kHz) dither line because of insufficient DAC range.
Attachment #4 puts everything into displacement noise units. The electronics noise of the coil driver / de-whitening circuit have not been included so at high frequencies, the projection is better than what will actually be realizable, but still well below the BHD requirement of 3e-17 m/rtHz.
|
Attachment 1: DACnoiseModels.pdf
|
|
Attachment 2: actuationRange.pdf
|
|
Attachment 3: deWhiteTFs.pdf
|
|
Attachment 4: dispNoiseModels.pdf
|
|
15782
|
Thu Jan 28 21:44:45 2021 |
gautam | Summary | BHD | HAM-A Coil Driver measurements After modifications |
Looks fine to me visually but the verdict can only be made once the z:p locations are quantitatively confirmed, and the noise tests pass. It would be interesting to see what kind of time-domain transient (in N of force) switching on the de-whitening introduces, i guess best done interferometrically.
Quote: |
I'll wait for comments until tomorrow to proceed with changes in the other board as well. I'll do noise measurements tomorrow.
|
|
15781
|
Thu Jan 28 18:04:55 2021 |
Anchal | Summary | BHD | HAM-A Coil Driver measurements After modifications |
I did the recommended modifications on of the boards with serial number S2100028. These included:
- R13, R27: 160 -> 75
- C11, C21: 470 nF -> 68nF
- C19: 4.7 uF -> 470 nF
- R15: 3.23 kOhm -> 1.82 kOhm
I took transfer function measurements with same method as in 40m/15774 and I'm presenting it here to ensure the modifications are correct and if I should proceed to the next board as well. I didn't have the data used to make plots in here but I think the poles and zeros have landed in the right spot. I'll wait for comments until tomorrow to proceed with changes in the other board as well. I'll do noise measurements tomorrow. |
Attachment 1: D1100117_S2100027_TF.pdf
|
|
Attachment 2: AfterChanges.zip
|
15780
|
Thu Jan 28 12:53:14 2021 |
Anchal | Summary | BHD | HAM-A Coil Driver measurements before modifications |
I took some steps to reduce the coupling of 60 Hz harmonics in noise measurement. The box was transferred to the floor instead of on top of another instrument. Measurement was immediately converted into single-ended using SR560 in battery mode with a gain of 10. All of the setups was covered in aluminum foil to increase isolation.
Spectrum measurement details
|
Attachment 1: D1100117_S2100027_Current_Noise_Spectrum.pdf
|
|
Attachment 2: D1100117_S2100027_Voltage_Noise_Spectrum.pdf
|
|
Attachment 3: D1100117_S2100028_Current_Noise_Spectrum.pdf
|
|
Attachment 4: D1100117_S2100028_Voltage_Noise_Spectrum.pdf
|
|
Attachment 5: SpectrumMeasurement.zip
|
15779
|
Tue Jan 26 15:37:25 2021 |
Anchal | HowTo | CDS | Acromag wiring investigation |
Here I present few wiring diagrams when using Acromag to avoid noisy behavior and ground loops.
Case 1: Only single-ended sources
- Attachment 1 gives a functioning wiring diagram when all sources are single ended.
- One should always short the RTN to IN- pin if the particular GND carried by that signal has not been shorted before to RTN for some other signal.
- So care is required to mark different GNDs of different powersupply separately and follow where they inadvertently get shorted, for example when a photodiode output is connected to FSS Box.
- Acromag should serve as the node of all GNDs concerned and all these grounds must not be connected to Earth GND at power supply ends or in any of the signal sources.
- I think this is a bit complicated thing to do.
Case 2: Some single and some differential sources
- Connect all single ended sources same as above keeping care of not building any ground loops.
- The differential source can be connected to IN+ and IN- pins, but there should be a high resistance path between IN- and RTN. See Attachment 2.
- Why this is the case, I'm not sure since I could not get access to internal wiring of Acromag (no response from them). But I have empirically found this.
- This helps IN- to float with respect to RTN according to the negative signal value. I've found that 10kOhm resistance works good. See 40m/15778.
- If RTN is shorted to nearby Earth GND (assuming none of the other power supply GNDs have been shorted to Earth GND) shows a reduction in noise for differential input. So this is recommended.
- This wiring diagram carries all complexity of previous case along with the fact that RTN and anything connected to it is at Earth GND now.
Case 3: Signal agnostic wiring
- Attachment 3 gives a wiring diagram which mimics the high resistance shorting of RTN to IN- in all ports regardless of the kind of signal it is used for reading.
- In this case, instead of being the node of the star configuration for GND, acromag gets detached from any ground loops.
- All differences in various GNDs would be kept by the power supplies driving small amounts of current through the 10 kOhm resistors.
- This is a much simpler wiring diagram as it avoids shorting various signal sources or their GNDs with each other, avoiding some of the ground loops.
Edit Wed Jan 27 13:38:19 2021 :
This solution is not acceptable as well. Even if it is successfull in reading the value, connecting resistor between IN- and RTN will not break the ground loops and the issue of ground loops will persist. Further, IN- connection to RTN breaks the symmetry between IN- and IN+, and hence reduces the common mode rejection which is the intended purpose of differential signal anyways. I'll work more on this to find a way to read differential signals without connecitng IN- and RTN. My first guess is that it would need the GND on the source end to be connected to EarthGND and RTN on acromag end to be connected to EarthGND as well.
|
Attachment 1: GeneralLabWiring.pdf
|
|
Attachment 2: GeneralLabWiring2.pdf
|
|
Attachment 3: GeneralLabWiring3.pdf
|
|
15778
|
Tue Jan 26 12:59:51 2021 |
Anchal | HowTo | CDS | Acromag wiring investigation |
Taking inspiration from SR785 on how it reads differential signal, I figured that acromag too always need a way to return current through RTN ports always. That must be the reason why everything goes haywire when RTN is not connected to IN-. Now for single ended signals, we can always short RTN to IN- and keep same GND but then we need to be careful in avoiding ground loops. I'm gonna post a wiring diagram in next post to show how if two signal sources connect to each other separately, a GND loop can be formed if we tie each IN- port to RTN on an acromag.
Coming to the issue of reading a differential signal, what SR785 does is that it connects 50 Ohm resistance between Earth GND and differential signal shields (which are supposed to signal GND). In a floating GND setting, SR785 connects a 1 MOhm resistor between input shield and Earth GND. This can be used to read a differential signal through a single BNC cable since the shiled can take arbitrary voltages thanks ti the 1 MOhm resistor.
We can do the same in acromag. Instead of shorting RTN to IN- ports, we can connect them through a large resistor which would let IN- float but will give a path for current to return through RTN ports. Attached here are few scenarios where I connected IN- to RTN throguh wire, 820 Ohms, 10kOhms and 1MOhms in two sub cases where RTN was left open or was shorted to Earth GND. In all cases, the signal was produced by a 9V battery outputing roughly 8.16V. It seems that 10kOhm resistor between RTN and IN- with RTN connected to Earth GND is the best scenario noise wise. I'll post more results and a wiring diagram soon. |
Attachment 1: TestingDifferentialSignalWithFloatingRTNwrtIN-.pdf
|
|
15777
|
Tue Jan 26 10:58:30 2021 |
gautam | Update | SUS | MC2 tickler stuck on |
For whatever reason, the autolocker didn't turn the tickle off for several hours. Seems to work okay now. The linked plot suggests that the coil balancing on MC2 is pretty lousy. |
15776
|
Mon Jan 25 18:18:04 2021 |
Anchal | Summary | BHD | Satellite Amplifier Transfer Functions and noise |
I took transfer function and noise measurement of satellite amplifier box's photodiode transimpedance circuit. For the measurement, I created a makeshift connector to convert backside DB25 into DB9 with the 4 channels for PDA input. The output was taken in differential form at the front PD Output port. To feed current to the circuit, I put in 12 kOhm resistors in series at the inputs, so the V/V transfer function measured was multiplied by 12 kOhm to get the transimpedance of the circuit.
Transfer Function Measurement details
- SR785 source out was fed into PDA input pins using a makeshift BNC-DB9-DB25 converter.
- The output from PDOut DB9 port was fed to test switch in D1900068 to separate differential signal.
- This differential signal was fed back to SR785 at input 2 in A-B configuration.
- Measurements are taken with file D1002818_TF.yml and D1002818_TF_LF.yml.
- A measurement of just cables without the DUT is taken as well.
- Commands.txt list all the commands used.
- All data is compiled and plotted in Plotting.ipynb
- D1100117_S2100029_TFandNoiseSpectrum.pdf shows all the transfer functions measured.
Spectrum Measurements
- Two pair of BNC cables were twisted together and clips were added at ends.
- One of the GND was connected to board GND. Rest were left unconnected to avoid ground loops.
- Each pair of signal was connected to PDOutP/N.
- The PDA inputs were shorted together to make zero input current to the board.
- Instrument noise with cables was measured by shorting the clips of the center cores and one of the shields of the two BNC cables together.
- Measurements were taken with file D1002818_SP.yml and D1002818_SP_LF.yml.
- Input referred current noise spectrum was calculated by dividing the output voltage noise spectrum by the measured transfer function.
- D1100117_S2100029_TFandNoiseSpectrum.pdf shows all the output votlage noise spectrum and input referred current noise spectrum measured.
Edit Wed Feb 10 15:14:13 2021 :
THE NOISE MEASUREMENT WAS WRONG HERE. SEE 40m/15799. |
Attachment 1: D1002818_S2100029_TFandNoiseSpectrum.pdf
|
|
Attachment 2: D1002818_Testing.zip
|
15775
|
Wed Jan 20 19:12:16 2021 |
gautam | Update | CDS | SLwebview fixed |
After fixing multiple issues, the model webviews are updating, should be done by tomorrow. It should be obvious from the timestamps on the index page which are the new ones. These new screens are better than the old ones and offer more info/details. Please look at them, and let me know if there are broken links etc. Once we are happy with this new webview, we can archive the old files and clean up the top directory a bit. I don't think this adds anything to the channel accounting effort but it's a nice thing to have up-to-date webviews, I found the LLO ones really useful in setting up the AS WFS model.
BTW, the crontab on megatron is set to run every day at 0844. The process of updating the models is pretty heavy because of whatever MATLAB overhead. Do we really need to have this run every day? I modified the crontab to run every other Saturday, and we can manually run the update when we modify a model. Considering this hasn't been working for ~3 years, I think this is fine, but if anyone has strong preference you can edit the crontab.
If someboy can volunteer to fix the MEDM screenshot that would be useful. |
15774
|
Wed Jan 20 18:07:09 2021 |
Anchal | Summary | BHD | HAM-A Coil Driver measurements before modifications |
I have taken transfer functions and noise measurements of the two HAM-A coil driver boxes D1100687 #S2100027 and #S2100028. All transfer functions look as expected. I'm not sure about the noise measurements. If anyone sees flaw in my measurement method, please let me know. I'm not sure why in some channels I got 10Hz harmoni peaks in the noise. That was very strange. Also let me know if my current noise estimate is wrong.
Transfer Function Measurement details
- SR785 source out was connected to the differential amplifier input of D1900068.
- The one pair of two BNC outputs of this differential amplifier goes directly to the SR785 Input 1 A and B.
- The DB9 output of the differential amplifier goes to the Coil Input DB9 connector J3.
- Header W2 was shorted to provide ground to the incoming signal.
- Header P4 was shorted to enable all the channels manually.
- Normal operation is the Acquisition mode (Acq) while when pins of header P3 are shorted, we go into the Run mode for respective channel.
- The “To Satellite Box” DB25 port at the read side was conencted to a DB25 breakout circuit and pins 1-9, 3-11, 5-13 and 7-15 were connected to 36 Ohm resistor to simulate Coil load.
- The “Output Monitor” on the rear side is then connected to the test switch DB9 port on D1900068.
- The the pair of BNCs from the test switch is connected to SR785 Input 2 A and B.
- Measurements are taken with file D1100687_TF.yml and D1100687_TF_LF.yml.
- A measurement of just cables without the DUT is taken as well.
- Commands.txt list all the commands used.
- All data is compiled and plotted in Plotting.ipynb
- D1100117_S2100027_TF.pdf and D1100117_S2100028_TF.pdf shows all the transfer functions measured.
Spectrum Measurements
- All channels were kept in disabled mode (Not shorting P4) to ensure their inputs are grounded on the board.
- I ran two BNC cables with their centers connected to output monitors V2+ and V2- and one of their shields connected to board GND.
- in SR785, A-B differential mode always runs with grounded shields mode, so effectively the board GND got grounded to SR785 GND through internal 50 Ohm resistor. But all ground loops have been evaded.
- The two BNC cables were twisted together to minimize the area between the two center cores of the cables as that is the remaining pickoff possible in this measurement.
- Instrument noise with cables was measured first but shorting the clips of the center cores and one of the shields of the two BNC cables together.
- Measurements were taken with file D1100687_SP.yml and D1100687_SP_LF.yml.
- D1100117_S2100027_Voltage_Noise_Spectrum.pdf and D1100117_S2100028_Voltage_Noise_Spectrum.pdf shows the measured voltage noise spectrum at the output monitors when loaded with 36 Ohms.
- D1100117_S2100027_Current_Noise_Spectrum.pdf and D1100117_S2100028_Current_Noise_Spectrum.pdf shows the esitmate current noise through the coil calculated by dividing the measured voltage noise by 2436 Ohms.
|
Attachment 1: MeasurementData.zip
|
Attachment 2: D1100117_S2100027_TF.pdf
|
|
Attachment 3: D1100117_S2100028_TF.pdf
|
|
Attachment 4: D1100117_S2100027_Voltage_Noise_Spectrum.pdf
|
|
Attachment 5: D1100117_S2100028_Voltage_Noise_Spectrum.pdf
|
|
Attachment 6: D1100117_S2100027_Current_Noise_Spectrum.pdf
|
|
Attachment 7: D1100117_S2100028_Current_Noise_Spectrum.pdf
|
|
15773
|
Wed Jan 20 10:13:06 2021 |
gautam | Update | Electronics | HV Power supply bypassing |
Summary:
Installing 10uF bypass capacitors on the High Voltage power supply line for the HV coil driver circuit doesn't improve the noise. The excess bump around a few hundred Hz is still present. How do we want to proceed?
Details
- The setup is schematically shown in Attachment #1.
- Physically, the capacitors were packaged into a box, see Attachment #2.
- This box is installed between the HVPS and the 2U chassis in which the circuit is housed, see Attachment #3.
- I measured the noise, (using the same setup as shown here to avoid exposing the SR785 input to any high voltage), for a variety of drive currents. To make a direct comparison, I took two sets of measurements today, one with the decoupling box installed and one without.
- The results are shown in Attachment #4. You can see there is barely any difference between the two cases. I've also plotted the expected noise per a model, and the measured Johnson noise of one of the 25kohm resistors being used (Ohmite, wirewound). I just stuck the two legs of the resistor into the SR785 and measured the differential voltage noise. There is a slight excess in the measured Johnson noise compared to what we would expect from the Fluctuation Dissipation theorem, not sure if this is something to be worried about or if it's just some measurement artefact.
Discussion:
So what do we do about this circuit? For the production version, I can make room on the PCB to install two 10uF film capacitors on the board itself, though that's unlikely to help. I think we've established that
- The excess noise is not due to the Acromag or the input Acromag noise filtering stage of the circuit, since the excess is present even when the input to the HV stage is isolated and shorted to ground.
- There was some evidence of coherence between the supply rails and the output of the HV stage (with input isolated and shorted to ground). The coherence had the "right shape" to explain the excess noise, but the maximum value was only ~0.1 (could have been because I was not measuring directly at the PA95's supply rail pins due to space constraints).
- The impedance of 10uF at 100Hz is ~150 ohms. idk what the impedance of the supply pins of the PA95 are at this frequency (this will determine the coupling of ripples in the HVPS output to the PA95 itself.
Do we have any better bipolar HV supply that I can use to see if that makes any difference? I don't want to use the WFS supplies as it's not very convenient for testing.
Not really related directly to this work but since we have been talking about current requirements, I attach the output of the current determining script as Attachment #5. For the most part, having 220ohm resistances on the new HAM-A coil driver boards will lead to ~half the DAC range being eaten up for the slow alignment bias. For things like MC1/MC3, this is fine. But for PRM/SRM/BS, we may need to use 100ohms. Chub has ordered all manner of resistances so we should have plenty of choices to pick from. |
Attachment 1: bypassCaps.pdf
|
|
Attachment 2: IMG_9079.jpg
|
|
Attachment 3: IMG_9078.jpg
|
|
Attachment 4: HVampNoise_driven_chassis.pdf
|
|
Attachment 5: printCoilCurrents.pdf
|
|
15772
|
Tue Jan 19 15:43:24 2021 |
gautam | Configuration | CDS | Updated CDS upgrade plan |
Not sure if 1Y1 can accommodate both c1sus2 and c1bhd as well as the various electronics chassis that will have to be installed. There may need to be some distribution between 1Y1 and 1Y3. Does Koji's new wiring also specify which racks hold which chassis?
Some minor improvements to the diagram:
- The GPS receiver in 1X7 should be added. All the timing in the lab is synced to the 1pps from this.
- We should add hyperlinks to the various parts datasheets (e.g. Dolphin switch, RFM switch, etc etc) so that the diagram will be truly informative and self-contained.
- Megatron and nodus, but especially chiara (NFS server), should be added to the diagram.
|
15771
|
Tue Jan 19 14:05:25 2021 |
Jon | Configuration | CDS | Updated CDS upgrade plan |
I've produced updated diagrams of the CDS layout, taking the comments in 15476 into account. I've also converted the 40m's diagrams from Omnigraffle ($150/license) to the free, cloud-based platform draw.io. I had never heard of draw.io, but I found that it has most all the same functionality. It also integrates nicely with Google Drive.
Attachment 1: The planned CDS upgrade (2 new FEs, fully replace RFM network with Gen 1 Dolphin IPC)
Attachment 2: The current 40m CDS topology
The most up-to-date diagrams are hosted at the following links:
Please send me any further corrections or omissions. Anyone logged in with LIGO.ORG credentials can also directly edit the diagrams. |
Attachment 1: 40m_CDS_Network_-_Planned.pdf
|
|
Attachment 2: 40m_CDS_Network_-_Current.pdf
|
|
15770
|
Tue Jan 19 13:19:24 2021 |
Jon | Update | CDS | Expansion chassis from LHO |
Indeed T1800302 is the document I was alluding to, but I completely missed the statement about >3 GHz speed. There is an option for 3.4 GHz processors on the X10SRi-F board, but back in 2019 I chose against it because it would double the cost of the systems. At the time I thought I had saved us $5k. Hopefully we can get the LLO machines in the near term---but if not, I wonder if it's worth testing one of these to see whether the performance is tolerable.
Can you please provide a link to this "list of boards"? The only document I can find is T1800302....
|
I confirm that PCIe 2.0 motherboards are backwards compatible with PCIe 1.x cards, so there's no hardware issue. My main concern is whether the obsolete Dolphin drivers (requiring linux kernel <=3.x) will work on a new system, albeit one running Debian 8. The OSS PCIe card is automatically configured by the BIOS, so no external drivers are required for that one.
Please also confirm that there are no conflicts w.r.t. the generation of PCIe slots, and the interfaces (Dolphin, OSSI) we are planning to use - the new machines we have are "PCIe 2.0" (though i have no idea if this is the same as Gen 2).
|
|
15769
|
Sat Jan 16 18:59:44 2021 |
gautam | Update | LSC | Modulation depth measurement |
I decided to analyze the data I took in December more carefully to see if there are any clues about the weird LSC sensing.
Attachment #1 shows the measurement setup.
- The PSL shutter was closed. All feedback to both lasers was disconnected during the measurement. I also disabled the input switch to the FSS Box - so the two laser beams being interfered shouldn't have any modulations on them other than the free running NPRO noise and the main IFO modulations.
- Everything is done in fiber as I had the beams already coupled into collimators and this avoided having to optimize any mode matching on the beat photodiode.
- The pickoff of the PSL is from the collimator placed after the triply resonant EOM that was installed for the air BHD experiment.
- The other beam is the EX laser beam, arriving at the PSL table via the 40m long fiber from the end (this is the usual beam used for ALS).
- I didn't characterize precisely the PLL loop shape. But basically, I wasn't able to increase the SR560 gain any more without breaking the PLL lock. Past experience suggests that the UGF is ~20 kHz, and I was able to get several averages on the AG4395 without the lock being disturbed.
Attachment #2 shows the measured spectrum with the PSL and EX laser frequency offset locked via PLL.
- The various peaks are identified.
- There are several peaks which I cannot explain - any hypothesis for what these might be? Some kind of Sorensen pollution? They aren't any multiples of any of the standard RF sources. They are also rather prominent (and stationary during the measurement time, which I think rules out the cause being some leakage light from the EY beam, which I had also left connected to the BeatMouth during the measurement).
- In the previous such characterization done by Koji, such spurious extra peaks aren't seen.
- Also, I can't really explain why some multiples of the main modulation are missing (could also be that my peak finding missed the tiny peaks)?
- The measuremet setup is very similar to what he had - important differences are
- Much of the optical path was fiber coupled.
- Beat photodiode is NF1611, which is higher BW than the PDA10CF.
- The second laser source was the Innolight EX NPRO as opposed to the Lightwave that was used.
- The RF source has been modified, so relative phasing between 11 MHz and 55 MHz is different.
Fitting the measured sideband powers (up to n=7, taking the average of the measured upper and lower sideband powers to compute a least squares fit if both are measured, else just that of the one sideband measured) agains those expected from a model, I get the following best fit parameters:

To be explicit, the residual at each datapoint was calculated as
.
The numbers compare favourably with what Koji reported I think - the modulation depths are slightly increased, consistent with the RF power out of the RF box being slightly increased after I removed various attenuators etc. Note the large uncertainty on the relative phase between the two modulations - I think this is because there are relatively few sidebands (one example is n=3) which has a functional dependence that informs on phi - most of the others do not directly give us any information about this parameter (since we are just measuring powers, not the actual phase of the electric field).
Attachment #3 shows a plot of the measured modulation profile, along with the expected heights plugging the best fit parameters into the model. The size of the datapoint markers is illustrative only - the dependence on the model parameters is complicated and the full covariance would need to be taken into account to put error bars on those markers, which I didn't do.
Attachment #4 shows a time domain measurement of the relative phasing between the 11 MHz and 55 MHz signals at the EOM drive outputs on the RF source box. I fit a model there and get a value for the relative phase that is totally inconsistent from what I get with this fit. |
Attachment 1: PLL.pdf
|
|
Attachment 2: modDepth.pdf
|
|
Attachment 3: modProfile.pdf
|
|
Attachment 4: EOMpath_postMod.pdf
|
|
15768
|
Fri Jan 15 17:04:45 2021 |
gautam | Update | LSC | Messed up LSC sensing |
I want to lock the PRFPMI again (to commission AS WFS). Have had some success - but in doing characterization, I find that the REFL port sensing is completely messed up compared to what I had before. Specifically, MICH and PRCL DoFs have no separation in either the 1f or 3f photodiodes.
- A sensing line driven in PRCL doesn't show up in the AS55 photodiode signal - this is good and as expected.
- For MICH - I set the MICH--->PRM actuation matrix element so as to minimize the height of the peak at the MICH drive frequency that shows up at the PRCL error point. My memory is that I used to be able to pretty much null this signal in the past, but I can't find a DTT spectrum in the elog as evidence. Anyways, the best effort nulling I can achieve now still results in a large peak at the PRCL error point. Since the sensing matrix doesn't actually make any sense, idk if it is meaningful to even try and calibrate the above qualitative statement into quantitative numbers of cross coupling in meters.
- With the PRMI locked on 1f error signals (ETMs misaligned, PRCL sensed with REFL11_I, MICH sensed with AS55_Q) - I tried tweaking the digital demod phase of the REFL33 and REFL165 signals. But I find that the MICH and PRCL peaks move in unison as I tweak the demod phase. This suggests to me that both signals are arriving optically in phase at the photodiode, which is weird.
- The phenomenon is seen also in the REFL11 signal.
I did make considerable changes to the RF source box, and so now the relative phase between the 11 MHz and 55 MHz signals is changed compared to what it was before. But do we really expect any effect even in the 1f signal? I am not able to reproduce this effect in simulation (Finesse), though I'm using a simplified model. I attach two sensing matrices to illustrate what i mean:
- Attachment #1 is in the PRFPMI state, with the IFO on RF control (CARM on REFL11, PRCL on REFL165_I, MICH on REFL165_Q, DARM on AS55_Q).
- Attachment #2 is between the transition to RF control (CARM and DARM on ALS, PRCL on REFL165_I, MICH on REFL165_Q). The CARM offset is ~4nm (c.f. the linewidth of ~20pm), so the circulating power in the arm cavities is low.
|
Attachment 1: PRFPMI_Jan12sensMat.pdf
|
|
Attachment 2: PRMI3f_ALS_Jan11_largeOffsetsensMat.pdf
|
|
15767
|
Fri Jan 15 16:54:57 2021 |
gautam | Update | CDS | Expansion chassis from LHO |
Can you please provide a link to this "list of boards"? The only document I can find is T1800302. In that, under "Basic Requirements" (before considering specific motherboards), it is specified that the processor should be clocked @ >3GHz. The 3 new supermicros we have are clocked at 1.7 GHz. X10SRi-F boards are used according to that doc, but the processor is clocked at 3.6 or 3.2 GHz.
Please also confirm that there are no conflicts w.r.t. the generation of PCIe slots, and the interfaces (Dolphin, OSSI) we are planning to use - the new machines we have are "PCIe 2.0" (though i have no idea if this is the same as Gen 2).
Quote: |
The motherboard actually has six PCIe slots and is on the CDS list of boards known to be compatible.
|
As for the CX4 cable - I still think it's good to have these on hand. Not good to be in a situation later where FE and expansion chassis have to be in different racks, and the copper cable can't be used. |
Attachment 1: Screenshot_2021-01-15_17-00-06.png
|
|
15766
|
Fri Jan 15 15:06:42 2021 |
Jon | Update | CDS | Expansion chassis from LHO |
Koji asked me assemble a detailed breakdown of the parts received from LHO, which I do based on the high-res photos that Gautam posted of the shipment.
Parts in hand:
Qty |
Part |
Note(s) |
2 |
Chassis body |
|
2 |
Power board and cooling fans |
As noted in 15763, these have the standard LIGO +24V input connector which we may want to change |
2 |
IO interface backplane |
|
2 |
PCIe backplane |
|
2 |
Chassis-side OSS PCIe x4 card |
|
2 |
CX4 fiber cables |
These were not requested and are not needed |
Parts still needed:
Qty |
Part |
Note(s) |
2 |
Host-side OSS PCIe x4 card |
These were requested but missing from the LHO shipment |
2 |
Timing slave |
These were not originally requested, but we have recently learned they will be replaced at LHO soon |
Issue with PCIe slots in new FEs
Also, I looked into the mix-up regarding the number of PCIe slots in the new Supermicro servers. The motherboard actually has six PCIe slots and is on the CDS list of boards known to be compatible. The mistake (mine) was in selecting a low-profile (1U) chassis that only exposes one of these slots. But at least it's not a fundamental limitation.
One option is to install an external PCIe expansion chassis that would be rack-mounted right above the FE. It is automatically configured by the system BIOS, so doesn't require any special drivers. It also supports hot-swapping of PCIe cards. There are also cheap ribbon-cable riser cards that would allow more cards to be connected for testing, although this is not as great for permanent mounting.
It may still be better to use the machines offered by Keith Thorne from LLO, as they're more powerful anyway. But if there is going to be an extended delay before those can be received, we should be able to use the machines we already have in conjunction with one of these PCIe expansion options. |
15765
|
Thu Jan 14 12:32:28 2021 |
gautam | Update | CDS | Rogue master may be doing something good? |
I think the "Rogue Master" setting on the RFM network may be doing some good. 5 mins, ago, all the CDS indicators were green, but I noticed an amber light on the c1rfm screen just now (amber = warning). Seems like at GPS time 1294691182, there was some kind of error on the RFM network. But the network hasn't gone down. I can clear the amber flag by running the global diag reset. Nevertheless, the upgrade of all RT systems to Dolphin should not be de-prioritized I think. |
Attachment 1: Screenshot_2021-01-14_12-35-52.png
|
|
15764
|
Thu Jan 14 12:19:43 2021 |
Jon | Update | CDS | Expansion chassis from LHO |
That's fine, we didn't actually request those. We bought and already have in hand new PCIe x4 cables for the chassis-host connection. They're 3 m copper cables, which was based on the assumption of the time that host and chassis would be installed in the same rack.
Quote: |
- Regarding the fibers - one of the fibers is pre-2012. These are known to fail (according to Rolf). One of the two that LHO shipped is from 2012 (judging by S/N, I can't find an online lookup for the serial number), the other is 2011. IIRC, Rolf offered us some fibers so we may want to take him up on that. We may also be able to use copper cables if the distances b/w server and expansion chassis are short.
|
|
15763
|
Thu Jan 14 11:46:20 2021 |
gautam | Update | CDS | Expansion chassis from LHO |
I picked the boxes up this morning. The inventory per Fil's email looks accurate. Some comments:
- They shipped the chassis and mounting parts (we should still get rails to mount these on, they're pretty heavy to just be supported on 4 rack nuts on the front). idk if we still need the two empty chassis that were requested from Rich.
- Regarding the fibers - one of the fibers is pre-2012. These are known to fail (according to Rolf). One of the two that LHO shipped is from 2012 (judging by S/N, I can't find an online lookup for the serial number), the other is 2011. IIRC, Rolf offered us some fibers so we may want to take him up on that. We may also be able to use copper cables if the distances b/w server and expansion chassis are short.
- The units are fitted with a +24V DC input power connector and not the AC power supplies that we have on all the rest of the chassis. This is probably just gonna be a matter of convenience, whether we want to stick to this scheme or revert to the AC input adaptor we have on all the other units. idk what the current draw will be from the Sorensen - I tested that the boards get power, and with noi ADCs/DACs/BIOs, the chassis draws ~1A (read off from DCPS display, not measured with a DMM). ~Half of this is for the cooling fans It seems like KT @ LLO has offered to ship AC power supplies so maybe we want to take them up on that offer.
- Without the host side OSSI PCIe card, timing interface board, and supermicro servers that actually have enough PCIe slots, we still can't actually run any meaningful test. I ran just a basic diagnostic that the chassis can be powered on, and the indicator LEDs and cooling fans run.
- Some photos of the contents are here. The units are stored along the east arm pending installation.
> Koji,
>
> Barebones on this order.
>
> 1. Main PCIe board
> 2. Backplane (Interface board)
> 3. Power Board
> 4. Fiber (One Stop) Interface Card, chassis side only
> 5. Two One Stop Fibers
> 6. No Timing Interface
> 7. No Binary Cards.
> 8. No ADC or DAC cards
>
> Fil Clara
>
|
15762
|
Wed Jan 13 16:09:29 2021 |
Anchal | HowTo | CDS | Acromag wiring investigation |
I'm working on a better wiring diagram that takes into account multiple power supplies, how their GND is passed forward to the circuits or sensors using those power supplies and what possible wiring configurations on Acromag would give low noise. I think I have two configurations in mind which I will test and update here with data and better diagrams.
I took some striptool images earlier yesterday. So I'm dumping them here for further comments or inferences. |
Attachment 1: SimpleTestsStriptoolImages.pdf
|
|
15761
|
Tue Jan 12 11:42:38 2021 |
gautam | HowTo | CDS | Acromag wiring investigation |
Thanks for the systematic effort.
- Can you please post some time domain plots (ndscope perferably or StripTool) to clearly show the different failure modes?
- The majority of our AI channels are "Referenced Single Ended Source" in your terminology. At least on the c1psl Acromag crate, there are no channels that are truly differential drive (case #3 in your terminology). I think the point is that we saw noisy inputs when the IN- wasn't connected to RTN. e.g. the thorlabs photodiode has a BNC output with the shield connected to GND and the central conductor carrying a signal, and that channel was noisy when the RTN was unconnected. Is that consistent with your findings?
- What is the prescription when we have multiple power supplies (mixture of Sorensens in multiple racks, Thorlabs photodiodes and other devices powered by an AC/DC converter) involved?
- I'm still not entirely convinced of what the solution is, or that this is the whole picture. On 8 Jan, I disconnected (and then re-connected) the FSS RMTEMP sensor from the Acromag box, to check if the sensor output was noisy or if it was the Acromag. The problem surfaced on Dec 15, when I installed some new electronics in the rack (though none of them were connected to the Acromag directly, the only common point was the Sorensen DCPS. And between 8 Jan and today, the noise RMS has decreased back to the nominal level, without me doing anything to the grounding. How to reconcile this?
|
15760
|
Tue Jan 12 08:21:47 2021 |
anchal | HowTo | CDS | Acromag wiring investigation |
I used an Acromag XT1221 in CTN to play around with different wiring and see what works. Following are my findings:
Referenced Single Ended Source (Attachment 1):
- If the source signal is referenced single ended, i.e. the signal is only on the positive output and the negative side is tied to GND on the source side AND this GND is also shared by the power supply powering the Acromag, then no additional wiring is required.
- The GND common to the power supply and the source is not required to be Earth GND but if done so, it should be at one point only.
- RTN terminal on Acromag can be left floating or tied to IN- terminal.
Floating Single Ended Source (Attachment 2):
- If the source signal is floating single-ended i.e. the signal is only on the positive output and the negative output is a floating GND on the source, the the IN- should be connected to RTN.
- This is the case for handheld calibrators or battery powered devices.
- Note that there is no need to connect GND of power supply to the floating GND on the source.
Differential Source (Attachment 3):
- If the source is differential output i.e. the signal is on both the positive output and the negative output, then connect one of the RTN terminals on Acromag to Earth GND. It has to be Earth GND for this to work.
- Note that you can no longer tie the IN- of different signals to RTN as they are all carrying different negative output from the source.
- Earth GND at RTN gives acromag a stable voltage reference to measure against the signals coming in IN+ and IN-. And the most stable voltage reference is of course Earth GND.
Conclusion:
- We might have a mix of these three types of signals coming to a single Acromag box. In that case, we have to make sure we are not connecting the different IN- to each other (maybe through RTN) as the differential negative inputs carry signal, not a constant voltage value.
- In this case, I think it would be fine to always use differential signal wiring diagram with the RTN connected to Earth GND.
- There's no difference in software configuration for the two types of inputs, differential or single-ended.
- For cases in which we install the acromag box inside a rack mount chasis with an associated board (example: CTN/2248), it is ok and maybe the best to use the Attachment 1 wiring diagram.
Comments and suggestions are welcome.
Related elog posts:
40m/14841 40m/15134
Edit Tue Jan 26 12:44:19 2021 :
Note that the third wiring diagram mentioned actually does not work. It is an error in judgement. See 40m/15762 for seeing what happens during this. |
Attachment 1: SingleEndedNonFloatingWiring.pdf
|
|
Attachment 2: SingleEndedFloatingWiring.pdf
|
|
Attachment 3: DifferentialSignalWiring.pdf
|
|
15759
|
Mon Jan 11 19:10:10 2021 |
rana | Update | BHD | Monte Carlo loop coupling Simulations |
- looking better, but the CARM plot still looks weird.
- you should plot from 0.01 - 10,000 Hz
- All of the loops should have true integrators below 1 Hz.
- I don't think these loops are stable; the Bode plot is not a good way to check stability for LTI systems since you can be fooled by phase wrapping.
|
15758
|
Mon Jan 11 16:11:51 2021 |
Yehonathan | Update | BHD | Monte Carlo loop coupling Simulations |
I dived into the alog to make the OLTFs in the MC_controls example more realistic. I was mainly inspired by these entries:
https://alog.ligo-la.caltech.edu/aLOG/uploads/47116_20190708131007_carmolg_20190702.png
https://alog.ligo-wa.caltech.edu/aLOG/index.php?callRep=18742
https://alog.ligo-wa.caltech.edu/aLOG/index.php?callRep=20466
and Evan's and Dennis's Theses.
Attachment 1 shows the new OLTFs. I tried to make them go like 1/f around the UGF and fall as quickly as possible at higher frequencies. I didn't do more advanced stability checks.
I also noticed that imbalances and detunings in the MC simulation can change the plants significantly. Especially DARM, CARM, and sometimes PRCL. I added the option to fix some OLTFs throughout the simulation. At every iteration, the simulation computes the required control filter to fix the selected OLTFs such that it will match the OLTFs in the undetuned and balanced IFO.
|
Attachment 1: MC_LANCE_OLTFs.pdf
|
|
15756
|
Fri Jan 8 20:01:11 2021 |
gautam | Update | ALS | Noisy ALS |
I did this test today. The excess noise around 100 Hz doesn't show up in the green beat.
See Attachment #1. The setup was as usual:
- X-Arm cavity length stabilized to PSL frequency using the POX locking loop.
- EX laser frequency locked to the X-Arm cavity length using the AUX PDH loop.
- The "BEATX" channel records frequency fluctuations in the beat sensed on the IR beat photodiode, while the "BEATY" channel records frequency fluctuations in the beat sensed on the Green beat photodiode.
- Since the green beat frequency fluctuations are twice that of the IR beat frequency fluctuations, I scaled the former ASD by a factor of 0.5 so as to compare apples to apples.
- At low frequencies, the green beat is noisier, but that channel doesn't show the excess noise at mid frequencies you see in the IR beat. So the AUX PDH sensing noise is not to blame I think.
So, this excess appears to truly be excess phase noise on the fiber (though I have no idea what the actual mechanicsm could be or what changed between Aug and Oct 2020 that could explain it. Maybe the HEPA?
For this work, I had to spend some time aligning the two green beams onto the beat photodiode. During this time, I shuttered the PSL, disabled feedback via the FSS servo, turned the HEPA high, and kept the EX green locked to the arm so as to have a somewhat stable beat signal I could maximize. Everything has been returned to nominal settings now (obviously, since I locked the arms to get the data).
You may ask, why do we care. In terms of RMS frequency noise, it would appear that this excess shouldn't matter. But in all my trials so far, I've been unable to transition control of the arm cavity lengths from POX/POY to ALS. I suppose we could try using the green beat, but that has excess low frequency noise (which was the whole point of the fiber coupled setup).
Quote: |
How about resurrecting the PSL table green beat for the X arm to see if the non-fiber setup shows the same level of the freq noise (e.g. the PDH locking became super noisy due to misalignment etc).
|
|
Attachment 1: ALSX_IR_green.pdf
|
|
15755
|
Thu Jan 7 23:25:19 2021 |
Koji | Update | ALS | Noisy ALS |
If the sensing noise level of the end PDH degraded for some reason, it'd make the out of loop stability worse without making the end pdh error level degraded.
It's just speculation.
|